Let f (x) = 4x^2 + 9 and g(x) = x – 3
(a) Find the composite function ( f o g)(x) and simplify. Show work.
(b) Find ( f o g ) (−1) . Show work.

Answer by stanbon(75887)

Let f (x) = 4x^2 + 9 and g(x) = x – 3
(a) Find the composite function ( f o g)(x) and simplify. Show work.
fog(x) = f[g(x)] = f(x-3) = 4(x-3)^2+9 = 4(x^2-6x+9)+9 = 4x^2 - 24x + 45
--------------------------------------------------
(b) Find ( f o g ) (−1) . Show work.
f[g(-1)] = f(-1-3) = f(-4) = 4(-4)^2+9 = 64+9 = 73
---------------
Cheers,
Stan H.
-----------